OpenStudy (anonymous):

What is the correct simplification of the expression (-2x + 4y)(3x - 7y). I got 6x-16xy-28y, Is it wrong? I used box method to get my answer.

OpenStudy (krabsen):

your answer is wrong

OpenStudy (krabsen):

x times x = x^2

OpenStudy (jackellyn):

Try FOIL: first, outside inside last. example: (-a+b)(c-d) -ac + (-a)(-d) + bc + b(-d) -ac + ad + bc -bd combine like terms (which wont be possible with my example) and that will give you the answer @CrazyC246

OpenStudy (anonymous):

That's what I was thinking of doing. My teacher told that if i don't understand FOIL use the box method.

OpenStudy (jackellyn):

@CrazyC246 , I am not familiar with the box method but FOIL is pretty straight up, just gotta watch your signs and multiplication.

OpenStudy (anonymous):



OpenStudy (anonymous):

This is what I did.

OpenStudy (jackellyn):

@CrazyC246 , remember when multiplying two variables with exponents x has an exponent 1 you add the exponents x^1(x^1)=x^(1+1)=x^2

OpenStudy (krabsen):